A 32.5 Ω resistor is connected to a power supply. As a result there is a current of 14.2 mA flowing through the resistor. What is the power dissipated in the resistor in milli-watts (mW)?
Answer #1

Power = i^2 * R OR V^2/R i is current r is resistance here we know both the values

(14.2 * 10^-3)^2 *32.5= 6.55 * 10^-3 W =6.66mW